The market price of a security is $64. Its expected rate of return is 13%. The risk-free rate is 6%, and the market risk premium is 9%. What will the market price of the security be if its beta doubles (and all other variables remain unchanged)? Assume the stock is expected to pay a constant dividend in perpetuity.
Answer #1

Beta =(Expected Rate-Risk free rate)/(Market Risk Premium =(13%-6%)/9% =7/9
Dividend =Price *Required Rate =64*13% =8.32
New Beta =2*7/9 =14/9
New Required rate =Risk free rate + beta*Market Risk Premium =6%+14/9*9% =20%
New Stock Price=Dividend/Required rate =8.32/20% =41.60